OpenStudy (anonymous):

Find the indicated dot product. r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w

OpenStudy (irishboy123):

v.w = <vx,vy,vz>•<wx,wy,wz> = vx*wx +vy*wy + vz*wz

OpenStudy (anonymous):

<-27, 28, 56> <18, -36, 49> 18 31 these are my answer choices, Is A the correct answer??

OpenStudy (astrophysics):

\[<3, 4, 7> \cdot <6, -9, 7> = <3 \times 6 + 4 \times -9 + 7 \times 7>\]

OpenStudy (astrophysics):

Just as irish boy stated.

OpenStudy (anonymous):

So its <18+(-36)+49> simplified down to 31??

OpenStudy (astrophysics):

Yeah

OpenStudy (anonymous):

thanks! :)